OpenStudy (anonymous):

solve on 0<=x<2pi sinx-cosx=1

OpenStudy (solomonzelman):

tired, but... sinx-cosx=1 sinx=1+cosx sin^2x=1+cos^2x +2cos(x) sin^2x=sin^2x + 2cos(x) 2cos(x)=0 cos(x)=0

OpenStudy (solomonzelman):

except that your interval is [0 , 2π)

OpenStudy (anonymous):

thanks

OpenStudy (solomonzelman):

Anytime, if that helps you.